Terry is buying fruit for a party. Apples are on sale for $.99 for an 8 ounce pack. Pears are on sale for $2.99 for a 6-pear pack. Each pear weighs 6 ounces. Write an inequality and tell which is the better deal per ounce.
Answer by Theo(13342)

8 ounces of apples = .99

6 * 6 ounces of pears = 36 ounces of pears = 2.99

apples = .99 / 8 = .12375 per ounce.

pears = 2.99 / 36 = .08305... per ounce.

pears are cheaper per ounce than apples.

your inequality would be 2.99 / 36 < .99 / 8

this simplifies to .08305... < .12375